OpenStudy (anonymous):

if sec theta = x + 1/x , prove that sec theta * tan theta = 2x

OpenStudy (anonymous):



OpenStudy (anonymous):

draw a triangle, label hypotenuse \[x+\frac{1}{x}\] and "adjacent side"1, so that \[\sec(\theta)=1+\frac{1}{x}\] then solve for the other side using pythagoras and take ratio. i didn't actually do it so let me write it on paper first, but i am fairly certain the solution will pop out

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

before i make fool of myself, is it \[x+\frac{1}{x}\] or \[\frac{x+1}{x}\]?

OpenStudy (anonymous):

can i have the whole answer please??

OpenStudy (anonymous):

x + 1/x

OpenStudy (anonymous):

first one or second one?

OpenStudy (anonymous):

first one

OpenStudy (aravindg):

help meeeeeee

OpenStudy (anonymous):

plz wait aravind

OpenStudy (anonymous):

help me first

OpenStudy (anonymous):

i am afraid i do not get 2x

OpenStudy (anonymous):

wat do u get

OpenStudy (anonymous):

in fact i can easily check that it is wrong by picking a value of x

OpenStudy (anonymous):

i'll get back to you. will try working it out. thank you

OpenStudy (aravindg):

helppppppppppp

OpenStudy (anonymous):

pick x = 1, then x+1/x=1+1=0 and if \[\sec(\theta)=2\] \[\cos(\theta)=\frac{1}{2}\] and so \[\sin(\theta)=\frac{\sqrt{3}}{2}\] and \[\tan(\theta)=\sqrt{3}\] and finally \[\sec(\theta)\tan(\theta)=2\times \sqrt{3}\neq 2\times 1\] so it is false